# Instruction-Free Tuning of Large Vision Language Models for Medical Instruction Following

### 19 Mar 2026 v1

@@ Abstract
keywords:
Instruction-Free Tuning, Large Vision Language Models, Instruction Tuning
[…]
[1]organization=Department of Electrical and Computer Engineering, The University of British Columbia, city=Vancouver, state=BC V6T 1Z4, country=Canada
[…]
[2]organization=Department of Robotics and Mechatronics Engineering, Daegu Gyeongbuk Institute of Science and Technology (DGIST), city=Daegu, postcode=42988, country=Republic of Korea
[…]
[3]organization=Division of Intelligent Robot, Daegu Gyeongbuk Institute of Science and Technology (DGIST), city=Daegu, postcode=42988, country=Republic of Korea
[…]
[4]organization=Vector Institute, city=Toronto, state=ON M5G 0C6, country=Canada
[…]
[5]organization=Department of Computer Science and Engineering, Pohang University of Science and Technology (POSTECH), city=Pohang, postcode=37673, country=Republic of Korea
@@ 1 Introduction
Large vision language models (LVLMs) have demonstrated strong general-purpose capabilities across a wide range of visual language understanding and generation tasks [38, 23]. This success is largely attributable to large-scale visual instruction tuning [24, 26], where models are fine-tuned on datasets consisting of image-instruction-output triplets (Fig. 1(a)). Despite its effectiveness in enhancing adherence to diverse human instructions, visual instruction tuning is highly dependent on the scale and quality of the instruction dataset [37, 40]. Therefore, instruction-tuned LVLMs tend to perform poorly in domains with limited publicly available data, such as medicine.
To mitigate this limitation, several studies have developed instruction datasets for the medical domain and leveraged them to specialize LVLMs for medical applications [21, 4]. In particular, medical images are typically accompanied by detailed textual descriptions (e.g., radiology reports or captions from PubMed Central); these studies exploit this inherent property of medical image-text pairs to construct datasets for visual instruction tuning. Specifically, these approaches build upon earlier successes in general-purpose LVLMs [31] by adapting language-only large language models (LLMs) to transform image descriptions into conversational image-instruction-output triplets for large-scale instruction dataset construction. However, the frequent use of domain-specific abbreviations and terminology [28], along with the need for expert-level contextual understanding [17], poses a significant challenge for constructing datasets using LLMs. Meanwhile, radiology reports typically begin with a lesion description and conclude with a diagnostic impression [16], which leads to strong inter-sentence confounding dependencies within the constructed datasets. These characteristics lead to the overlooking of visual features and exacerbate hallucinations [9], thereby necessitating visual instruction tuning that accounts for these factors. Therefore, we revisit fine-tuning approaches that reduce reliance on handcrafted or generated instructions (as illustrated in Fig. 1(b)), aiming to avoid inefficiencies in constructing instruction datasets and to mitigate potential drawbacks caused by inadequate instructions.
@@
Instruction-free tuning refers to fine-tuning LVLMs using only image-description pairs (e.g., radiology reports or captions), without relying on any instructions. To the best of our knowledge, this is the first work to formalize address this task for instruction-tuned LVLMs. However, fine-tuning a model without human-curated instructions introduces additional challenges, particularly when it relies on the inherent properties of paired raw data as text instructions (e.g., “Describe this medical scan”), regardless of whether LLMs are used during dataset construction [19, 5]. In practice, paired descriptions (e.g., radiology reports or captions) do not perfectly align with the natural response (e.g., “The image…”) that would typically follow a conversational instruction (e.g., “Describe…”). The misalignment between the instruction-description pairs in the fine-tuning and pre-training datasets poses a risk of degrading the instruction-following capability of the pre-trained LVLM. Specifically, this may lead to image-dependence, where the model loses its pre-trained instruction-following capability and generates fixed-style responses that rely excessively on the image regardless of the given instruction (see Fig. 8). For instance, a model fine-tuned for report generation may fail to respond to other types of queries (e.g., “Is the texture typical…”) and struggle to generalize to question-answering tasks beyond report writing. Alternatively, this may result in instruction-dependence (or parameter-dependence), where the model responds appropriately only to instructions encountered during fine-tuning and fails to generate responses aligned with the fine-tuning data for unseen instructions (instead respond based on the pre-trained knowledge; see other LVLMs in Table 1). Furthermore, due to the nature of instruction-following LVLMs, instructions seen during fine-tuning are not guaranteed to be provided at inference time, highlighting the importance of preventing overfitting to such instructions.
@@
In response to these challenges, we introduce a proxy instruction that is optimized to align with the set of image-description pairs in the fine-tuning dataset. Unlike conventional fine-tuning, which uses text instructions and optimizes model parameters, our method jointly updates both learnable proxy instructions and model parameters. This enables the LVLM to adapt to domain-specific datasets without degrading its instruction-following capability by employing a well-aligned instruction. As a result, the fine-tuned model can flexibly respond to a wide range of text instructions in the medical context during inference. Furthermore, we refined the proxy instruction through an exponential moving average, i.e., a momentum proxy instruction. This mitigates overfitting to a specific proxy instruction and allows the remaining parameters to learn compensated representations that reflect overall fine-tuning trends. Meanwhile, fine-tuning a model while maintaining its pre-trained instruction-following capability can mitigate the challenges posed by the unique characteristics of medical data. Since the diagnostic conclusion can often be inferred from the earlier words alone (e.g., findings), the model is likely to largely overlook visual features when generating the latter part of the response. In this context, we propose a strategy, response shuffling, which randomly shuffles the sentence order of the ground truth description, thereby preventing the model from relying too heavily on previous words. Notably, employing a proxy instruction disregards the negative effects of shuffling output order during fine-tuning, enabling a simple strategy to substantially enhance medical visual question answering (VQA) performance.
@@ 4.2 Data Preparation for Instruction-Free Evaluation
When converting to multiple-choice VQA, attribute annotations were formatted into question-answer pairs, usually with five options. The question was constructed using the attribute name in the form “What is the {name}?”, whereas for the SKINCON dataset, the fixed question “What is the dermatologic feature?” was used. For multiple-choice options, the annotated attribute value was set as the correct answer, while randomly sampled non-overlapping values from the same attribute served as the incorrect options. Attributes with fewer than five options (e.g., yes/no) were converted into a limited choice set (e.g., two options). For MIMIC-CXR evaluation, we used the Medical-CXR-VQA dataset [14]. This dataset consists of VQA pairs for chest X-ray interpretation built from the MIMIC-CXR dataset, covering questions on abnormalities, presence, view, location, level, and type. Since we used only frontal chest X-rays for our experiments, we excluded view-related question types. For consistency with the other datasets, we added incorrect options to convert the dataset into a multiple-choice VQA format with five options. Incorrect options were set to different values within the same question type, while yes/no questions were limited to two options. For abnormality-type questions with multiple disease answers, we performed VQA multiple times, considering each abnormality as a separate correct choice. To promote reproducibility and fair comparison, we release the full instruction-free tuning benchmark for the SKINCON, WBCAtt, and CBIS datasets, including the preprocessing pipelines for the medical reports and the multiple-choice VQA evaluation protocols.
@@ 5.1.1 Main Results
5.1.2 Comparison with Instruction-Free Tuning Variants
@@ 5.3.2 Discussion of Misalignment
Figure 6: Ablation study on the momentum coefficient $\alpha$ on the SKINCON, WBCAtt, and CBIS datasets.
Figure 7: Ablation study on the number of instruction vectors NN on the SKINCON, WBCAtt, and CBIS datasets.
@@ 5.3.3 Ablation on Coefficients and Instruction Scale
Fig. 6 shows the accuracy of the ablation study on the momentum coefficient. The value 0.999 yielded the highest accuracy, indicating that excessively fast or slow adaptation may limit performance improvements, similar to [13]. Fig. 7 shows the accuracy of the ablation study on the number of instruction vectors. The highest accuracy was achieved with 8 instruction vectors, whereas increasing the number to 32 resulted in lower performance. This may be related to general questions averaging about 8 tokens in length.

### 24 Apr 2026 v2

@@ 1 Introduction
Large vision language models (LVLMs) have demonstrated strong general-purpose capabilities across a wide range of visual language understanding and generation tasks [18]. This success is largely attributable to large-scale visual instruction tuning, where models are fine-tuned on datasets consisting of image-instruction-output triplets (Fig. 1(a)). Despite its effectiveness in enhancing adherence to diverse human instructions, visual instruction tuning is highly dependent on the scale and quality of the instruction dataset [33]. Therefore, instruction-tuned general-purpose LVLMs tend to perform poorly in domains with limited publicly available data, such as medicine.
@@
Several studies have adapted LVLMs for medical applications by using instruction datasets specifically designed for the medical domain [26, 4]. Since medical images are typically paired with detailed textual descriptions (e.g., radiology reports or captions from PubMed Central), these studies leverage language-only large language models (LLMs) to transform image descriptions into conversational image-instruction-output triplets for visual instruction tuning. However, the frequent use of domain-specific abbreviations and terminology [24], along with the need for expert-level contextual understanding [16], poses a significant challenge for constructing datasets using LLMs. Another challenge is that, radiology reports typically begin with a lesion description and conclude with a diagnostic impression [15], which leads to strong inter-sentence confounding dependencies within the constructed datasets. These characteristics lead to the overlooking of visual features and exacerbate hallucinations [8], thereby necessitating visual instruction tuning that accounts for these factors. Therefore, we revisit fine-tuning approaches that reduce reliance on handcrafted or generated instructions (as illustrated in Fig. 1(b)), aiming to avoid inefficiencies in constructing instruction datasets and to mitigate potential drawbacks caused by inadequate instructions.
@@
Instruction-free tuning refers to fine-tuning LVLMs using only image-description pairs (e.g., radiology reports or captions), without relying on any instructions. To the best of our knowledge, this is the first work to formalize address this task for instruction-tuned LVLMs. However, fine-tuning a model without human-curated instructions introduces additional challenges, particularly when it leverages on the inherent properties of paired raw data as text instructions (e.g., “Describe this medical scan”) [4]. Paired descriptions (e.g., radiology reports or captions) do not perfectly align with the natural response (e.g., “The image…”) that would typically follow a conversational instruction (e.g., “Describe…”). The misalignment between the instruction-description pairs in the fine-tuning and pre-training datasets poses a risk of degrading the instruction-following capability of the pre-trained LVLM.
Specifically, this may lead to image-dependence, where the model loses its pre-trained instruction-following capability and generates fixed-style responses that rely excessively on the image regardless of the given instruction (see Fig. 7). For instance, a model fine-tuned for report generation may fail to respond to other types of queries (e.g., “Is the texture typical…”) and struggle to generalize to question-answering tasks beyond report writing. Additionally, this may result in instruction-dependence (or parameter-dependence), where the model responds appropriately only to instructions encountered during fine-tuning and fails to generate responses aligned with the fine-tuning data for unseen instructions (instead respond based on the pre-trained knowledge; see other LVLMs in Table 1). Furthermore, due to the nature of instruction-following LVLMs, instructions seen during fine-tuning are not guaranteed to be provided at inference time, highlighting the importance of preventing overfitting to such instructions.
@@
In response to these challenges, we introduce a proxy instruction that is optimized to align with the set of image-description pairs in the fine-tuning dataset. Unlike conventional fine-tuning, which uses text instructions and optimizes model parameters, our method jointly updates both learnable proxy instructions and model parameters. This enables the LVLM to adapt to domain-specific datasets without degrading its instruction-following capability through a well-aligned instruction. As a result, the fine-tuned model can flexibly respond to a wide range of text instructions in the medical context during inference. Furthermore, we refined the proxy instruction through an exponential moving average, i.e., a momentum proxy instruction. This mitigates overfitting to a specific proxy instruction and allows the remaining parameters to leverage compensated representations that reflect overall fine-tuning trends.
In addition, diagnostic conclusions can often be inferred from the preceding text alone (e.g., findings), leading models to overlook visual features when generating the later part of the response. In this context, we propose a strategy, response shuffling, which randomly shuffles the sentence order of the ground truth description, thereby preventing the model from relying too heavily on previous words. Notably, employing a proxy instruction disregards the negative effects of shuffling output order during fine-tuning, enabling a simple strategy to substantially enhance medical visual question answering (VQA) performance.
@@ 4.2 Data Preparation for Instruction-Free Evaluation
When converting to multiple-choice VQA, attribute annotations were formatted into question-answer pairs, usually with five options. The question was constructed using the attribute name in the form “What is the {name}?”, whereas for the SKINCON dataset, the fixed question “What is the dermatologic feature?” was used. For multiple-choice options, the annotated attribute value was set as the correct answer, while randomly sampled non-overlapping values from the same attribute served as the incorrect options. Attributes with fewer than five options (e.g., yes/no) were converted into a limited choice set (e.g., two options). For MIMIC-CXR evaluation, we used the Medical-CXR-VQA dataset [13]. This dataset consists of VQA pairs for chest X-ray interpretation built from the MIMIC-CXR dataset, covering questions on abnormalities, presence, view, location, level, and type. Since we used only frontal chest X-rays for our experiments, we excluded view-related question types. For consistency with the other datasets, we added incorrect options to convert the dataset into a multiple-choice VQA format with five options. Incorrect options were set to different values within the same question type, while yes/no questions were limited to two options. For abnormality-type questions with multiple disease answers, we conducted VQA separately for each disease, setting each abnormality as a separate correct choice. To promote reproducibility and fair comparison, we release the full instruction-free tuning benchmark for the SKINCON, WBCAtt, and CBIS datasets, including the preprocessing pipelines for the medical reports and the multiple-choice VQA evaluation protocols.
@@ 5.1.1 Main Results
5.1.2 Comparison with Instruction-Free Tuning Variants
@@ 5.3.2 Discussion of Misalignment
Figure 6: Ablation study of (a) the momentum coefficient $\alpha$ and (b) the number of instruction vectors NN on the SKINCON, WBCAtt, and CBIS datasets.
@@ 5.3.3 Ablation on Coefficients and Instruction Scale
Fig. 6(a) shows the accuracy of the ablation study on the momentum coefficient. The value 0.999 yielded the highest accuracy, indicating that excessively fast or slow adaptation may limit performance improvements, similar to [12]. Fig. 6(b) shows the accuracy of the ablation study on the number of instruction vectors. The highest accuracy was achieved with 8 instruction vectors, whereas increasing the number to 32 resulted in lower performance. This may be related to general questions averaging about 8 tokens in length.